A rope, 12 meter long, is cut into two pieces at a random point between its ends. What
is the probability that the length of one piece is at least thrice the length of the other?
1
Expert's answer
2021-08-15T18:04:03-0400

Let "X=" the length of the shorter piece of rope, so that "X\leq 6\ m."

The random variable "X" is uniformly distributed on the interval "0\leq X\leq 6," so


"P(0\leq X\leq x)=\dfrac{x}{6}"

If "\dfrac{X}{12-X}\leq\dfrac{1}{3}," so


"12-X\geq3X"

"X\leq3"

"P(0\leq X\leq 3)=\dfrac{3}{6}=\dfrac{1}{2}"

The probability that the length of one piece is at least thrice the length of the other is "\dfrac{1}{2}."